# supplement

### v1

@@ 7 Conclusion
Via the CDF transformation, we consider a suite of functional time-series models to model and forecast subnational age distribution of death counts for 47 Japanese prefectures. The CDF transformation is ideal to handle the zero values in the subnational data. Within the CDF transformation, we also present two general strategies for constructing pointwise prediction intervals.
By dividing the datasets into training and testing samples, we evaluate and compare point and interval forecast accuracy among the functional time-series models. We confirm the importance of joint modeling techniques, in which the MFTS, MLFTS, FANOVA, and HDFPCA models generally perform well on the aggregate level across 47 prefectures. We present a novel heatmap to show the frequency (of 47 prefectures), in which one method performs the best for each forecast horizon. The individual forecast errors for various horizons, obtained from all methods for each prefecture, are available in a developed shiny app at https://cfjimenezv.shinyapps.io/Shiny_App_CDF/.
There are several ways the methodology can be further extended, and we briefly list four. First, the functional standard deviation in (8) was computed coordinate-wise. A range of functional depth exists, which can be implemented to compute other variants of standard deviations. Second, instead of symmetric prediction intervals, asymmetric ones can be considered, searching for two tuning parameters to adjust the lower and upper bounds. Third, the data set was divided equally into the training, validation, and testing samples. Other proportions may be possible and lead to a more accurate selection of tuning parameter $\xi_{\alpha}$ and more accurate interval forecasts. Fourth, to demonstrate we implemented a suite of functional time-series models. Other time-series extrapolation models may also be considered, such as the robust functional principal component analysis by Oguamalam et al. (2025). It extends Mahalanobis distance to Bayes spaces and introduces a regularized approach for improved principal component estimation in the presence of outliers.
@@
    Code for shiny application.The [Uncaptioned image] code to produce a shiny user interface for the point and interval forecasts for the 47 Japanese prefectures. The [Uncaptioned image] codes are available at the following repository: https://github.com/cfjimenezv07/HDFTS_CDF/tree/main/Shiny_App_CDF
@@ Acknowledgment
This research was supported by the Australian Research Council Discovery Project (grant no. DP230102250) and Future Fellowship (grant no. FT240100338). The author is grateful to Dr Yang Yang from the University of Newcastle for assistance with the heatmaps.
@@
In Table LABEL:tab:appendix_1, averaged across 47 prefectures, we compare the point forecast accuracy among the functional time-series models, when the number of retained components is set to be six as in Hyndman et al. (2013). From the summary statistics of the KLD and JSD, the MFTS offers the most accurate point forecasts for the female and male data.
Table 3: Averaged across 47 prefectures, we evaluate and compare the point forecast accuracy, measured by the KLD and JSD, among the functional time-series models. The number of components is determined by setting $K=6$.
@@
KLD 1 0.005 0.006 0.005 0.005 0.004 0.004 0.004 0.004
2 0.007 0.007 0.006 0.006 0.005 0.005 0.004 0.005
3 0.008 0.008 0.007 0.008 0.006 0.005 0.005 0.005
4 0.010 0.010 0.009 0.009 0.007 0.006 0.006 0.006
5 0.013 0.012 0.011 0.011 0.008 0.007 0.007 0.007
6 0.015 0.013 0.012 0.013 0.009 0.008 0.008 0.008
7 0.018 0.015 0.014 0.016 0.010 0.009 0.009 0.009
8 0.023 0.017 0.017 0.019 0.011 0.010 0.010 0.010
9 0.029 0.021 0.021 0.024 0.012 0.011 0.011 0.012
10 0.035 0.025 0.025 0.029 0.013 0.013 0.012 0.013
11 0.041 0.029 0.028 0.034 0.014 0.014 0.011 0.013
12 0.048 0.026 0.033 0.039 0.015 0.010 0.012 0.012
13 0.057 0.032 0.042 0.048 0.016 0.011 0.014 0.014
14 0.070 0.040 0.049 0.057 0.018 0.013 0.016 0.015
15 0.096 0.050 0.068 0.079 0.019 0.014 0.020 0.021
16 0.128 0.067 0.097 0.108 0.020 0.016 0.029 0.028
Mean 0.038
0.024
0.028 0.032 0.012
0.010
0.011 0.011
Median 0.026
0.019
0.019
0.022 0.012
0.010
0.011 0.011
JSD 1 0.037 0.038 0.036 0.036 0.033 0.031 0.030 0.031
2 0.040 0.041 0.039 0.039 0.036 0.033 0.033 0.034
3 0.044 0.044 0.042 0.042 0.038 0.035 0.035 0.036
4 0.049 0.047 0.046 0.047 0.041 0.038 0.038 0.038
5 0.054 0.051 0.050 0.051 0.044 0.040 0.040 0.041
6 0.059 0.054 0.054 0.055 0.047 0.042 0.043 0.043
7 0.065 0.057 0.058 0.060 0.050 0.044 0.045 0.045
8 0.072 0.060 0.063 0.065 0.054 0.047 0.048 0.048
9 0.079 0.065 0.069 0.072 0.056 0.049 0.050 0.050
10 0.088 0.071 0.076 0.080 0.059 0.052 0.053 0.053
11 0.097 0.076 0.083 0.088 0.061 0.054 0.054 0.054
12 0.108 0.078 0.091 0.097 0.063 0.052 0.056 0.056
13 0.120 0.086 0.101 0.107 0.065 0.055 0.060 0.059
14 0.133 0.093 0.110 0.117 0.068 0.059 0.064 0.063
15 0.157 0.104 0.131 0.140 0.070 0.061 0.071 0.072
16 0.186 0.123 0.158 0.168 0.073 0.064 0.085 0.084
Mean 0.087
0.068
0.076 0.079 0.054
0.047
0.050 0.050
Median 0.075
0.063
0.066 0.068 0.055
0.048
0.049 0.049
@@
In Table LABEL:tab:appendix_2, averaged across 47 prefectures, we compare the interval forecast accuracy among the functional time-series models. From the summary statistics of the CPD and interval scores at the level of significance $\alpha=0.2$, the MFTS offers the smallest CPD value for the female data, while the MLFTS offers the smallest CPD values for the male data. In terms of the interval scores, the MFTS provides the smallest interval scores for the female data, while the MFTS and MLFTS present about the same and equally smallest interval scores for the male data.
@@
CPD 1 0.054 0.040 0.046 0.056 0.052 0.043 0.043 0.049
2 0.064 0.042 0.051 0.060 0.054 0.051 0.043 0.050
3 0.068 0.045 0.059 0.068 0.056 0.056 0.047 0.054
4 0.076 0.049 0.064 0.083 0.056 0.065 0.049 0.058
5 0.083 0.051 0.069 0.092 0.061 0.073 0.056 0.066
6 0.086 0.052 0.071 0.099 0.059 0.076 0.052 0.073
7 0.087 0.053 0.072 0.105 0.059 0.079 0.052 0.067
8 0.091 0.056 0.071 0.112 0.063 0.081 0.055 0.074
9 0.096 0.056 0.075 0.111 0.073 0.080 0.059 0.078
10 0.104 0.056 0.082 0.125 0.073 0.080 0.061 0.081
11 0.103 0.050 0.078 0.122 0.074 0.085 0.068 0.085
12 0.106 0.058 0.082 0.113 0.065 0.084 0.067 0.076
13 0.106 0.065 0.084 0.113 0.067 0.078 0.069 0.078
14 0.116 0.068 0.091 0.122 0.066 0.077 0.069 0.074
15 0.121 0.068 0.086 0.122 0.056 0.069 0.064 0.065
Mean 0.091
0.054
0.072 0.100 0.062 0.072
0.057
0.068
Median 0.091
0.053
0.072 0.111 0.061 0.077
0.056
0.073
$\overline{S}_{0.2}$ 1 232 244 227 242 242 219 204 229
2 278 275 265 294 271 247 230 254
3 324 309 301 348 302 275 257 281
4 388 347 349 428 333 303 287 309
5 454 392 396 504 370 334 315 341
6 525 438 435 587 402 361 341 368
7 599 477 486 678 431 378 367 393
8 693 526 552 785 454 401 393 418
9 798 587 625 915 482 422 421 444
10 925 673 716 1068 513 453 453 478
11 1041 747 799 1210 529 482 448 495
12 1174 764 902 1364 559 452 465 512
13 1329 869 1042 1557 574 484 514 560
14 1474 956 1159 1753 644 540 590 633
15 1863 1145 1445 2124 865 698 819 868
Mean 806
583
647 924 465
403
407 439
Median 693
526
552 785 454 401
393
418
@@
In machine learning, a popular methodology known as conformal prediction (Shafer & Vovk, 2008) is used to construct probabilistic forecasts calibrated on out-of-sample errors. From the absolute value of $[\widehat{\varepsilon}_{1,s}^{g}(u),\dots,\widehat{\varepsilon}_{M,s}^{g}(u)]$, we compute its $100(1-\alpha)\%$ quantile for a level of significance $\alpha$, denoted by $q_{\alpha}(u)$. The prediction interval can be obtained as
@@
The conformal prediction works well when the sample size in the validation set is reasonably large (Dhillon et al., 2024). For a small set of samples, we suggest the standard deviation approach in Section 5. In Table LABEL:tab:appendix_3, the MFTS provides the smallest CPD values under the EVR criterion, while the MLFTS offers the smallest CPD values under $K=6$ for the female data. For the male data, the MLFTS is recommended. In terms of interval scores, the MFTS and MLFTS perform on par and are equally the best for the female data. For the male data, the MLFTS is preferable.
Table 5: Averaged across 47 prefectures, we evaluate and compare the interval forecast accuracy, measured by the CPD and interval score, among the functional time-series models at the level of significance $\alpha=0.2$. The number of components is determined by the EVR criterion or $K=6$; for each we highlight the method with the smallest overall CPD and interval score values.
@@
CPD F 1 0.080 0.096 0.055 0.096 0.098 0.067 0.051 0.047 0.067
2 0.087 0.087 0.054 0.110 0.049 0.075 0.045 0.052 0.079
3 0.093 0.082 0.058 0.124 0.105 0.083 0.052 0.057 0.091
4 0.105 0.078 0.062 0.151 0.043 0.096 0.054 0.064 0.110
5 0.116 0.077 0.066 0.161 0.128 0.108 0.059 0.067 0.123
6 0.124 0.073 0.063 0.172 0.047 0.118 0.061 0.065 0.136
7 0.129 0.066 0.066 0.170 0.122 0.123 0.058 0.068 0.140
8 0.133 0.063 0.066 0.172 0.056 0.130 0.058 0.068 0.145
9 0.138 0.063 0.071 0.177 0.121 0.136 0.063 0.072 0.156
10 0.147 0.074 0.082 0.186 0.049 0.147 0.075 0.083 0.171
11 0.151 0.083 0.082 0.192 0.138 0.150 0.086 0.091 0.172
12 0.148 0.085 0.090 0.186 0.052 0.150 0.090 0.099 0.166
13 0.151 0.093 0.099 0.193 0.151 0.150 0.098 0.107 0.167
14 0.171 0.121 0.124 0.212 0.212 0.169 0.122 0.132 0.185
15 0.214 0.159 0.165 0.263 0.263 0.208 0.165 0.173 0.222
Mean 0.132 0.087
0.080
0.171 0.098 0.127
0.076
0.083 0.142
Median 0.133 0.082
0.066
0.172 0.098 0.130
0.061
0.068 0.145
M 1 0.076 0.109 0.056 0.094 0.083 0.066 0.059 0.055 0.058
2 0.078 0.112 0.063 0.104 0.096 0.073 0.069 0.058 0.066
3 0.078 0.114 0.069 0.109 0.101 0.074 0.076 0.065 0.072
4 0.082 0.124 0.086 0.121 0.119 0.079 0.087 0.074 0.082
5 0.088 0.131 0.095 0.129 0.124 0.090 0.098 0.086 0.099
6 0.098 0.140 0.105 0.134 0.137 0.102 0.114 0.096 0.115
7 0.096 0.131 0.096 0.125 0.134 0.106 0.113 0.091 0.119
8 0.099 0.130 0.098 0.125 0.127 0.110 0.113 0.093 0.126
9 0.107 0.132 0.095 0.127 0.134 0.120 0.120 0.091 0.134
10 0.114 0.140 0.098 0.128 0.153 0.125 0.133 0.099 0.137
11 0.106 0.147 0.095 0.122 0.164 0.121 0.139 0.096 0.141
12 0.108 0.142 0.092 0.115 0.162 0.118 0.140 0.099 0.139
13 0.107 0.142 0.090 0.106 0.162 0.119 0.142 0.096 0.138
14 0.117 0.163 0.097 0.124 0.172 0.135 0.160 0.106 0.156
15 0.144 0.192 0.142 0.149 0.196 0.161 0.192 0.142 0.181
Mean 0.100 0.137
0.092
0.121 0.138 0.107 0.117
0.090
0.127
Median 0.099 0.132
0.095
0.124 0.134 0.110 0.114
0.093
0.116
$\overline{S}_{0.2}$ F 1 255 386 241 279 323 235 249 230 248
2 302 412 277 329 329 283 282 269 301
3 351 442 312 387 379 331 318 306 358
4 421 473 358 476 352 400 356 356 441
5 489 504 407 553 467 468 400 408 517
6 565 526 446 642 377 543 439 451 603
7 642 554 494 734 527 618 477 502 695
8 741 581 551 846 432 714 516 563 807
9 853 628 620 979 609 826 570 639 941
10 986 686 699 1147 456 961 638 725 1104
11 1106 756 766 1287 745 1084 705 801 1249
12 1242 734 859 1443 467 1220 713 904 1415
13 1383 822 994 1639 899 1361 817 1035 1608
14 1565 909 1098 1832 490 1541 920 1152 1802
15 1887 1018 1321 2184 1181 1852 1054 1395 2142
Mean 853 629 630 984
535
829 564 649 949
Median 741 581 551 846
467
714 516 563 807
M 1 258 273 218 266 247 233 220 205 229
2 282 297 241 287 275 260 247 230 253
3 306 323 262 301 299 287 274 256 278
4 329 350 289 321 331 313 301 285 302
5 356 373 311 342 360 344 330 311 331
6 377 391 329 354 385 369 353 334 355
7 397 407 346 368 408 394 372 356 377
8 418 425 369 384 429 415 391 380 404
9 444 443 389 403 463 444 411 404 429
10 462 470 414 422 513 462 443 428 455
11 484 495 408 440 561 484 468 424 479
12 490 453 422 443 596 493 436 436 490
13 509 482 471 474 660 509 467 486 525
14 560 524 521 510 716 555 509 533 567
15 618 560 625 586 874 609 555 624 631
Mean 419 418
374
393 475 411 385
379
407
Median 418 425
369
348 429 415 391
380
404

### v2

@@ 7 Conclusion
Via the CDF transformation, we consider a suite of functional time-series models to model and forecast the subnational age distribution of death counts for 47 Japanese prefectures. The CDF transformation is ideal for handling zero values in subnational data. Within the CDF transformation, we also present two general strategies for constructing pointwise prediction intervals.
By dividing the datasets into training and testing samples, we evaluate and compare the point and interval forecast accuracy among the functional time-series models. We confirm the importance of joint modeling techniques, in which the MFTS, MLFTS, FANOVA, and HDFPCA models generally perform well on the aggregate level across 47 prefectures. We present a novel heatmap showing the frequency (of 47 prefectures), in which one method performs the best for each forecast horizon. Individual forecast errors for various horizons, obtained from all methods for each prefecture, are available in a developed shiny app at https://cfjimenezv.shinyapps.io/Shiny_App_CDF/. The forecast life-table death counts are not only important to compute age-specific survival probability and life expectancy in demography, but they are also useful to determine annuity prices for various ages and maturities studied in Shang & Haberman (2020).
There are several ways in which the methodology can be extended, and we briefly list five. 1) The functional standard deviation in (8) was computed coordinate-wise. There is a range of functional depth that can be implemented to compute other variants of standard deviations. 2) Instead of symmetric prediction intervals, asymmetric ones can be considered, searching for two tuning parameters to adjust the lower and upper bounds. 3) Third, the data set was divided almost equally into training, validation, and testing samples. Other proportions may be possible, leading to a more accurate selection of the tuning parameter $\xi_{\alpha}$ and more accurate interval forecasts. 4) We implemented a suite of functional time-series models for evaluation and comparison. Other time-series extrapolation models may also be considered, such as the robust functional principal component analysis by Oguamalam et al. (2025). It extends the Mahalanobis distance to Bayes spaces and introduces a regularized approach for improved principal component estimation in the presence of outliers. 5) We focus on modeling and forecasting the conditional mean. For modeling and forecasting conditional heteroskedasticity, the functional generalized autoregressive conditional heteroskedasticity (GARCH) model may be more adequate (see, e.g., Aue et al., 2017; Cerovecki et al., 2019; Rice et al., 2020 and 2023; Kearney et al., 2025). Future work will be to incorporate the GARCH framework into some functional models considered.
@@
    Code for a shiny application.The [Uncaptioned image] code to produce a shiny user interface for the point and interval forecasts for the 47 Japanese prefectures. The [Uncaptioned image] codes are available at the following repository: https://github.com/cfjimenezv07/HDFTS_CDF/tree/main/Shiny_App_CDF
@@ Acknowledgment
This research was supported by the Australian Research Council Discovery Project (grant no. DP230102250) and Future Fellowship (grant no. FT240100338). We are grateful for the many insightful comments and suggestions from the seminar participants at the University of Melbourne and the University of New South Wales. The author is grateful to Dr Yang Yang from the University of Newcastle for assistance with the heatmaps.
@@
In Table LABEL:tab:appendix_1, averaged across 47 prefectures, we compare the point forecast accuracy among the functional time-series models, when the number of retained components is set to six as in Hyndman et al. (2013). Based on the summary statistics of the KLD and JSD, the MFTS provides the most accurate point forecasts for the female and male data.
Table 4: Averaged across 47 prefectures, we evaluate and compare the point forecast accuracy of the functional time-series models, measured by the KLD and JSD. The number of components is determined by setting $K=6$.
@@
KLD 1 0.004 0.004 0.004 0.006 0.003 0.003 0.003 0.004
2 0.006 0.008 0.006 0.007 0.004 0.004 0.004 0.005
3 0.008 0.009 0.008 0.008 0.005 0.005 0.005 0.005
4 0.010 0.011 0.009 0.009 0.006 0.006 0.006 0.006
5 0.012 0.013 0.010 0.011 0.006 0.007 0.006 0.007
6 0.014 0.014 0.011 0.013 0.007 0.007 0.007 0.007
7 0.017 0.016 0.013 0.015 0.008 0.008 0.008 0.008
8 0.021 0.018 0.016 0.018 0.009 0.009 0.009 0.009
9 0.026 0.021 0.019 0.022 0.010 0.010 0.010 0.011
10 0.032 0.025 0.023 0.027 0.011 0.012 0.011 0.011
11 0.039 0.029 0.027 0.033 0.012 0.013 0.013 0.013
12 0.047 0.035 0.034 0.041 0.012 0.015 0.014 0.014
13 0.058 0.042 0.040 0.048 0.013 0.017 0.015 0.015
14 0.067 0.038 0.047 0.063 0.014 0.011 0.015 0.018
15 0.088 0.053 0.067 0.083 0.014 0.014 0.020 0.023
16 0.116 0.071 0.089 0.110 0.015 0.018 0.027 0.030
17 0.151 0.085 0.106 0.125 0.015 0.020 0.031 0.033
Mean 0.042 0.029 0.031 0.038 0.010 0.011 0.012 0.013
Median 0.026 0.021 0.019 0.022 0.010 0.010 0.010 0.010
JSD 1 0.032 0.034 0.033 0.037 0.027 0.026 0.026 0.031
2 0.040 0.042 0.039 0.040 0.033 0.032 0.032 0.034
3 0.044 0.046 0.043 0.043 0.035 0.034 0.034 0.036
4 0.048 0.049 0.047 0.047 0.038 0.037 0.037 0.038
5 0.052 0.053 0.050 0.051 0.040 0.039 0.039 0.040
6 0.057 0.056 0.053 0.055 0.043 0.041 0.041 0.042
7 0.062 0.059 0.057 0.059 0.045 0.043 0.043 0.044
8 0.067 0.062 0.061 0.065 0.048 0.045 0.045 0.047
9 0.074 0.066 0.066 0.070 0.051 0.047 0.048 0.049
10 0.081 0.070 0.072 0.077 0.053 0.049 0.050 0.051
11 0.090 0.076 0.079 0.087 0.055 0.051 0.052 0.054
12 0.101 0.083 0.088 0.098 0.057 0.053 0.055 0.057
13 0.115 0.090 0.098 0.109 0.059 0.056 0.058 0.060
14 0.128 0.094 0.109 0.124 0.060 0.055 0.061 0.066
15 0.149 0.111 0.130 0.144 0.061 0.060 0.069 0.074
16 0.175 0.128 0.150 0.169 0.063 0.066 0.079 0.085
17 0.201 0.137 0.164 0.181 0.063 0.069 0.085 0.089
Mean 0.089 0.074 0.079 0.086 0.049 0.047 0.050 0.053
Median 0.074 0.066 0.066 0.070 0.051 0.047 0.048 0.049
@@
In Table LABEL:tab:appendix_2_K6, averaged across 47 prefectures, we compare the interval forecast accuracy among the functional time-series models. From the summary statistics of the CPD and interval scores at the level of significance $\alpha=0.2$, the MFTS offers the smallest CPD value for the female data, while the MLFTS offers the smallest CPD values for the male data. In terms of interval scores, the MFTS yields the smallest for the female data, while FANOVA yields the smallest for the male data.
@@
CPD 1 0.052 0.042 0.045 0.058 0.054 0.049 0.046 0.051
2 0.059 0.038 0.046 0.061 0.056 0.058 0.050 0.055
3 0.066 0.041 0.057 0.068 0.055 0.060 0.052 0.057
4 0.073 0.045 0.061 0.079 0.058 0.068 0.056 0.061
5 0.072 0.045 0.059 0.084 0.058 0.070 0.054 0.062
6 0.080 0.052 0.065 0.095 0.061 0.073 0.058 0.071
7 0.084 0.049 0.066 0.097 0.059 0.073 0.054 0.069
8 0.089 0.052 0.068 0.105 0.061 0.078 0.055 0.071
9 0.090 0.054 0.073 0.106 0.070 0.077 0.060 0.067
10 0.101 0.055 0.082 0.115 0.071 0.077 0.068 0.078
11 0.108 0.056 0.087 0.128 0.075 0.084 0.076 0.118
12 0.109 0.054 0.087 0.127 0.074 0.081 0.071 0.129
13 0.113 0.054 0.076 0.127 0.071 0.075 0.075 0.128
14 0.119 0.064 0.085 0.129 0.067 0.076 0.081 0.155
15 0.120 0.091 0.093 0.135 0.065 0.080 0.084 0.156
16 0.109 0.103 0.104 0.140 0.062 0.062 0.088 0.153
Mean 0.089 0.056 0.073 0.112 0.065 0.071 0.063 0.094
Median 0.090 0.053 0.067 0.108 0.062 0.073 0.058 0.071
$\overline{S}_{0.2}$ 1 233 253 234 248 239 221 208 225
2 281 292 275 301 268 252 237 251
3 326 325 310 352 294 278 261 275
4 374 357 346 406 324 305 285 299
5 421 388 376 460 348 327 304 321
6 485 433 421 533 381 354 330 350
7 558 480 468 614 403 372 353 369
8 649 531 533 721 428 392 380 395
9 734 587 602 825 450 408 404 418
10 845 659 689 961 472 436 429 438
11 969 756 800 1117 486 460 455 457
12 1134 848 896 1305 513 485 471 487
13 1283 872 999 1507 516 451 489 513
14 1508 1035 1220 1804 552 504 572 573
15 1834 1226 1466 2122 615 578 714 656
16 2130 1463 1754 2410 840 797 942 903
Mean 885 667 841 1376 484 465 502 460
Median 734 587 689 1305 472 436 455 457
@@
In machine learning, a popular methodology known as conformal prediction (Shafer & Vovk, 2008) is used to construct probabilistic forecasts calibrated on out-of-sample errors. From the absolute value of $\widehat{\varepsilon}_{1,s}^{g}(u),\dots,\widehat{\varepsilon}_{M,s}^{g}(u)]$, we compute its $100(1-\alpha)\%$ quantile for a level of significance $\alpha$, denoted by $q_{\alpha}(u)$. The prediction interval can be obtained as
@@
The conformal prediction approach performs well when the validation set has a reasonably large sample size (Dhillon et al., 2024). For smaller sample sets, we recommend using the standard deviation method described in Section 5. As shown in Table LABEL:tab:appendix_3, the MLFTS achieves the lowest CPD values under the EVR criterion, whereas the MFTS provides the smallest CPD values for $K=6$ in the female dataset. For the male dataset, MLFTS is the preferred method. Regarding interval scores, MFTS attains the lowest values under the EVR criterion, while HDFPCA performs best for $K=6$ in the female data. For males, UFTS and FANOVA are favored under the EVR criterion, with MFTS performing best for $K=6$.
Table 6: Averaged across 47 prefectures, we evaluate and compare the interval forecast accuracy, measured by the CPD and interval score, for the functional time-series models at the significance level $\alpha=0.2$. The number of components is determined by the EVR criterion or $K=6$; for each, we highlight the method with the smallest overall CPD and interval score values.
@@
CPD F 1 0.090 0.109 0.063 0.090 0.102 0.067 0.056 0.054 0.067
2 0.105 0.106 0.065 0.105 0.052 0.080 0.057 0.056 0.080
3 0.120 0.099 0.066 0.120 0.118 0.091 0.057 0.062 0.091
4 0.136 0.094 0.067 0.136 0.044 0.103 0.059 0.066 0.103
5 0.141 0.084 0.062 0.141 0.130 0.109 0.060 0.064 0.109
6 0.159 0.083 0.063 0.159 0.052 0.124 0.062 0.068 0.124
7 0.161 0.080 0.071 0.161 0.133 0.133 0.064 0.073 0.133
8 0.163 0.074 0.071 0.163 0.060 0.138 0.063 0.071 0.138
9 0.172 0.077 0.075 0.172 0.130 0.147 0.068 0.077 0.147
10 0.183 0.080 0.081 0.183 0.054 0.161 0.075 0.084 0.161
11 0.192 0.081 0.096 0.192 0.135 0.171 0.084 0.095 0.171
12 0.199 0.095 0.106 0.199 0.055 0.178 0.095 0.108 0.178
13 0.198 0.095 0.107 0.198 0.168 0.176 0.099 0.112 0.176
14 0.221 0.115 0.129 0.221 0.082 0.188 0.115 0.132 0.188
15 0.256 0.158 0.162 0.256 0.241 0.219 0.158 0.164 0.219
16 0.299 0.190 0.186 0.299 0.138 0.256 0.191 0.201 0.256
Mean 0.184 0.103 0.096 0.184 0.117 0.152 0.092 0.101 0.152
Median 0.177 0.094 0.078 0.177 0.118 0.136 0.071 0.081 0.136
M 1 0.094 0.122 0.062 0.094 0.090 0.058 0.064 0.063 0.058
2 0.113 0.127 0.072 0.113 0.111 0.070 0.077 0.068 0.070
3 0.119 0.130 0.076 0.119 0.115 0.076 0.083 0.070 0.076
4 0.125 0.137 0.083 0.125 0.117 0.083 0.091 0.077 0.083
5 0.123 0.137 0.087 0.123 0.122 0.090 0.098 0.080 0.090
6 0.133 0.146 0.097 0.133 0.137 0.105 0.109 0.090 0.105
7 0.134 0.139 0.099 0.134 0.142 0.111 0.107 0.090 0.111
8 0.126 0.132 0.095 0.126 0.139 0.117 0.112 0.089 0.117
9 0.116 0.127 0.093 0.116 0.139 0.118 0.110 0.087 0.118
10 0.119 0.133 0.096 0.119 0.154 0.119 0.117 0.092 0.119
11 0.114 0.129 0.100 0.114 0.160 0.118 0.118 0.099 0.118
12 0.110 0.130 0.102 0.110 0.177 0.122 0.118 0.103 0.122
13 0.100 0.121 0.093 0.100 0.172 0.121 0.114 0.091 0.121
14 0.104 0.130 0.105 0.104 0.187 0.125 0.125 0.099 0.125
15 0.115 0.143 0.118 0.115 0.196 0.133 0.136 0.114 0.133
16 0.134 0.183 0.151 0.134 0.234 0.156 0.179 0.151 0.156
Mean 0.120 0.135 0.097 0.120 0.142 0.104 0.107 0.092 0.104
Median 0.119 0.133 0.096 0.119 0.138 0.118 0.113 0.091 0.118
$\overline{S}_{0.2}$ F 1 286 423 247 286 331 253 256 237 253
2 338 459 285 338 323 308 297 280 308
3 393 489 319 393 411 360 331 316 360
4 457 521 354 457 349 419 366 353 419
5 514 548 386 514 470 474 398 386 474
6 593 573 429 593 393 551 438 436 551
7 671 601 477 671 549 632 481 486 632
8 784 627 536 784 451 741 521 549 741
9 897 660 600 897 624 855 570 617 855
10 1050 715 676 1050 477 1001 633 707 1001
11 1217 778 766 1217 735 1170 710 804 1170
12 1396 877 862 1396 513 1356 802 908 1356
13 1588 865 954 1588 968 1563 820 1003 1563
14 1872 1015 1181 1872 570 1843 994 1237 1843
15 2185 1199 1387 2185 1368 2153 1201 1469 2153
16 2528 1343 1552 2528 619 2475 1358 1665 2475
Mean 1092 731 740 1092 603 1095 630 742 1095
Median 840 644 638 840 495 928 546 677 928
M 1 259 280 221 259 251 224 222 209 224
2 281 309 247 281 286 249 253 238 249
3 296 335 268 296 314 272 278 262 272
4 312 359 287 312 334 291 303 284 291
5 320 378 301 320 353 311 325 302 311
6 337 399 319 337 383 337 350 325 337
7 351 412 337 351 413 361 367 347 361
8 366 424 359 366 438 385 383 372 385
9 379 435 378 379 462 404 396 390 404
10 394 457 397 394 504 420 420 412 420
11 414 475 426 414 548 440 444 439 440
12 440 507 445 440 626 468 472 456 468
13 456 454 454 456 699 485 432 464 485
14 500 498 545 500 839 528 481 551 528
15 562 555 661 562 977 586 543 657 586
16 609 600 744 609 1094 640 595 730 640
Mean 393 419 396 393 516 395 387 414 395
Median 372 418 378 372 450 400 375 381 400
Appendix D: Model Confidence Set (MCS) Analysis
We performed a rigorous statistical comparison of the point forecast accuracy across all competing models using the Model Confidence Set (MCS) procedure (Hansen et al., 2011). The MCS procedure identifies a set of superior models, $\mathcal{M}^{*}$, where the null hypothesis of Equal Predictive Ability (EPA) cannot be rejected at a given significance level $\alpha$.
The analysis was based on the point forecast errors measured by the Kullback-Leibler Divergence (KLD) for nine Functional Time-Series (FTS) models over 17 forecast horizons ($h=1$ to $h=17$). The procedure was applied to the forecast errors for 47 Japanese prefectures separately for male and female mortality data. We used a significance level of $\alpha=0.10$, a block-bootstrap approach with $B=2000$ resamples, and the $T_{\max}$ statistic.
The heatmaps in Figure 9 visualize the results of the MCS procedure. In each cell, the count represents the number of prefectures (out of 47) for which the corresponding model was included in the statistically superior set $\mathcal{M}^{*}$ at that horizon, demonstrating the robust predictive ability of the models across the regional series. For the female data, the MLFTS method consistently achieved statistical superiority across most prefectures at shorter horizons. Conversely, the HDFPCA method proved most robust at longer horizons. For the male dataset, the MFTS method was consistently included in the superior set $\mathcal{M}^{*}$ for both short and long forecast horizons.
(a) Female data
(b) Male data
Figure 9: Heatmaps summarizing the results of the Model Confidence Set (MCS) procedure. The analysis compares the point forecast accuracy, measured by the Kullback-Leibler Divergence (KLD), across nine functional time-series models over 17 forecast horizons ($h=1$ to $h=17$) for both female and male mortality data. Each cell value indicates the number of the 47 prefectures for which the corresponding model was included in the superior set, $\mathcal{M}^{*}$, at that specific horizon (each row sums to 47). Cells are color-coded based on the count: Green highlights models included in the superior set for six or more prefectures ($\geq 6$), Red for four or fewer prefectures ($\leq 4$), and White for exactly five.